OpenStudy (anonymous):

A 30-kg child rides on a circus Ferris wheel that takes her around a vertical circular path with a radius of 20 m every 22 s. What is the magnitude of the resultant force on the child at the highest point on this trajectory?

OpenStudy (anonymous):

Note: it is given that the period of the rotation is 22s. From this: \(v=\frac{2 \pi r}{T}\) Then you have the velocity. Now, note also that the centripetal force, is actually the resultant force! By \(F_c = m \frac{v^2}{r}\)

OpenStudy (anonymous):

This is a ferris wheel and the forces on the child is Normal force upward and force of tension downward. So the equation is Force of tension - Normal force = FC. Is it not right?

OpenStudy (anonymous):

In a ferris wheel, the tension does not act on the child...right? She isn't hung by a steel cable... The resultant force, or \(\Sigma F=F_c\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm assuming, that the girl doesn't just stick to the wheel. But you should get the answer that way. Good luck! :)